Chapter 18, Problem 14 GO The dranings show three charges that have the same magnitude but may haive different signs. In all cases the distance d between the charges is the same. The magnitude of the charges is lo-7.8 C, and the distance between them is ฮ«J 8 mm. Determine the magnitude of the net force on charge 2 foreach of the tr ee e ing. (o) F- (b) F- (c) F-
Answer #1

Given q 7.811C 7.8 * 10-0C

d-3.8mm-0,0038m

(a)q2 is positive .q1 is negative,so pulling q2 towards left and q3 is pushing q2 towards left

Add the Forces

F_{1}=kq_{1}q^{2}/d^{2}

Fi-9 * 109 * 7.8 * 10-oC * 7.8 * 10-6C/(0.0038m)

Fi-9 10 7.8 10 7.8100.0038

F1 0.54756/0.0038-2

F1-379 19.67N

since charges are identical and distance is same

F3-Fi = 37919·67N

Net Force F = F1 + F3 = 2F, = 2 * 379 19.67N

ANSWER: F75839.34N

============================

b)

all charges are positive, so forces are repulsive

q1 is pushing q2 towards right and q3 is pushing q2 towards left

since the forces are equal and opposite,they cancel out

Net force is zero

ANSWER: {color{Red} F=0N}

============================

c)

q2 is negative .q1 is positive ,so pulling q2 towards left

q3 is positive ,so pulling q2 downwards

From (a)F1-379 19.67N (towards left)

Fa = 379 19.67M (downwards)

Net Force ,F=sqrt{F_{1}^{2}+F_{3}^{2}}

F V/37919.672 379 19.672

ANSWER: F 53626.51

========================